Having issues solving these 2 problems:
1) How many kilograms of antifreeze must be added to 4 kilograms of a 10% antifreeze solution to produce a 20 % solution?
2) How many grams of water must be evaporated from 10 grams of a 40% antiseptic solution to produce a 50% solution.

Answer by mananth(16946)

-------------- percent ---------------- Amount
Antifreeze solution 10 ---------------- 4 kg
Antifreeze 100 ---------------- x kg
Mixture 20 ---------------- 4 + x kg

10*4+100x=20(4+ x)
40+100 x=80+20 x
100x-20 x=-40+ 80
80 x = 40
/ 80
x = 0.5 kg Antifreeze

100 grams solution contains 60 gms water
600=50(10+x)
600=500+50x
100=50x
x=2 gms of water has to be removed
m.ananth@hotmail.ca